• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name P. Monaghan

    Log entry time 00:18:57 on March14,2005

    Entry number 141411

    keyword=Swing Shift Summary

    Swing shift Summary

    Shift Workers:
    Peter Monaghan
    Elena Kuchina

    Kinematics: Physics production with 40uA beam on the tilted carbon target.

    16:00 Shift begins with run 2631 in progress; beam delivery reasonably stable.

    17:30 MCC called to inform us that they are having vacuum problems in the injector (again! - they had these yesterday too!) and beam will be off for half an hour to an hour.

    17:50 Beam back - eralier than expected too!

    18:50 MCC called to tell us that the 'injector-on-call' person had arrived and was going to investigate their vacuum problems in the injector. Trip rate for the last hour was high because the vacuum in the injector is fluctuating and causing the current to be lost. MCC thinks we should be down for about thirty minutes.

    19:30 MCC are now ready to restore beam back into the hall.

    20:15 MCC are still encountering problems with delivering a stable beam - two injector experts have now been called in and are currently working on the situation. The beam has been tripping off almost constantly in the last twenty-five minutes, so MCC are now only able to deliver a reduced current of about 17uA at the moment in the hope that it will be stable.

    20:25 MCC take beam away, probably for about 30 minutes to continue working on the injector problems.

    21:17 Beam back - MCC delivering 40uA again.

    21:24 MCC take beam away again - still having problems and are trying to fix them; maybe another 30 minutes.

    22:15 After talking to the program deputy and discussing with the run coordinator, we are going to take 20uA beam instead of 40uA. In order to deliver the higher current, MCC would have to make an access into the tunnel to do some repairs/test/troubleshooting etc with the necessary experts on site as well. We decided to run tonight with whatever MCC can reliably deliver; hopefully, the problems will be fixed tomorrow.

    22:30 MCC called to tell us that they are going to have problems even at 20uA and so they have decided to take the beam away to go into the tunnel and switch 'guns'. They estimate this is going to take 2-3 hours - d'oh!

    23:50 Brad and Peter made a controlled access to try and fix the communication problems with the BigBite high voltage crates. Alas, no success!

    Run Summary

    15:51 - 16:27 2631 200K
    16:27 - 17:06 2632 200K
    17:07 - 17:28 2633 83K Run ended early due to MCC problems.
    17:50 - 18:36 2634 200K
    18:36 - 18:50 2635 40K Run ended early due to MCC problems.
    19:33 - 20:25 2636 150K Beam tripping; MCC have more problems.
    21:17 - 21:24 2637 2K JUNK; MCC unable to deliver good beam.

    Now running with 20uA beam.

    Changed prescale factors,

    ps1=50; ps2=5; ps3=5; ps4=30; ps5=1; ps6=60000; ps7=60000; ps8=99999.

    22:20 - 22:30 2638 51K MCC took beam away to change guns.